OpenStudy (anonymous):

Calculate the slope, if defined, of the straight line through the given pair of points. Try to do the problem without writing anything down except the answer. HINT [See Quick Examples on page 79.] (If an answer is undefined, enter UNDEFINED.) (0, 0) and (1, 4)

OpenStudy (anonymous):

over one, up four, slope is four

OpenStudy (campbell_st):

to find the gradient \[m = (y _{2} - y _{1})/(x _{2} - x _{1})\] so m = (4-0)/(1-0) m = 4